OpenStudy (anonymous):

How many years would it take for an investment to triple at 6% APR compounded annually

OpenStudy (anonymous):

Annual compounding: \[ 3 = (1+.06)^n \]Where \(n\) is the number of years. So, working from this: \[ \log_{1.06}(3)=n\implies\\ \frac{\ln(3)}{\ln(1.06)}=n\approx 18.8542\text{ years} \]

OpenStudy (anonymous):

Thank you very much!!

OpenStudy (anonymous):

Sure thing.